OpenStudy (anonymous):

If given f(x) = x + 4 and g(x) = –2x2 + 3x – 1, what is (f – g)(x)? A. (f – g)(x) = –2x2 + 4x – 3 B. (f – g)(x) = –2x2 – 2x + 5 C. (f – g)(x) = 2x2 + 2x + 5 D. (f – g)(x) = 2x2 – 2x + 5

OpenStudy (anonymous):

f(x)-g(x)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[(f-g)(x) = f(x) - g(x) = x+4 - (-2x^2+3x-1)\]Carefully expand that. Don't forget that the - sign in front of the parentheses changes the sign of each term of the stuff inside the parentheses!